OpenStudy (anonymous):

If circumference of circle is 1.57 m what is radius?

OpenStudy (anonymous):

@mojojojo

TheSmartOne (thesmartone):

\[C=2\pi \times r\]

TheSmartOne (thesmartone):

\[1.57=2\pi \times r\]

TheSmartOne (thesmartone):

\[r=\frac{ 1.57 }{ 2\pi }\]

OpenStudy (anonymous):

ok that's what I was guessing thanks now I wont forget

TheSmartOne (thesmartone):

np

OpenStudy (anonymous):

it wont calculate right

OpenStudy (anonymous):

2.4649 that doesn't seem right

TheSmartOne (thesmartone):

\[2\pi=6.28318530718\]

OpenStudy (anonymous):

thanks

TheSmartOne (thesmartone):

\[\frac{ 1.57 }{ 6.28318530718}=0.24987326065\]

OpenStudy (anonymous):

thanks could u look at my other problem